The Eight-Fold way June 4, 2007

Physics 213 General Physics I

GALILEO GALILEI′S N OBLE EIGHT FOLD PAT H T O KIN EMAT ICAL N IRVAN A

1. Right Coordinates. Begin by drawing the situation. Only by including thy coordinate axes may thou discern the direction in which to proceed.

2. Right Instants. Meditate and identify each important instant in the problem. If necessary, return to the beginning and expand thy drawing to clearly indicate every one.

3. Right Variables. Name each of thy servants, the kinematical variables x, v, a, et cetera. If a variable be different at different instants, distinguish it by an appropriate subscript. Although thy servants are—as all things are—without self, one must be consistent in these labels; much of human suffering is caused by the mixing of names.

4. Right Quantities. With this cleansing performed, reread the problem. Does it ask for an instant or time interval? A position or displacement? An average or instantaneous quantity? With the knowledge gained in the first three steps, thou canst now identify the known quantities and their values. Never forget that some data may be implicit in the problem statement.

5. Right Equations. Meditate deeply on the kinematical equations. Which relate desired quantities to known quantities? Meditation is not linear. Multiple equations may be needed at a given instant, or the same equation may be needed multiple times at different instants.

6. Right Calculation. Thou shalt now enter a turbulent period. Only by solving symbolically for the desired quantity canst thou achieve true mindfulness.

7. Right Numerics. With mindfulness now reached, substitute in thy numbers and convert thy units.

8. Right Interpretation. Gaze upon thy final answer to determine if it is at all reasonable. If it is not, repeat thy meditation. Failure to do so may result in rebirth back into this course rather than the realization of Kinematical Nirvana.

Acknowledgement

Many thanks are due to Michael Buice for proofreading this document and also for contributing his knowelege of Buddhism.

Copyright Notice

This document is copyrighted © 2006-7 by Itai Seggev. It may, however, be freely redistributed or used in derivative works for noncommercial purposes and with proper attribution under the terms of the Creative Commons Attribution-NonCommercial-ShareAlike License, Version 3.0. This license should be held to apply equally to all forms of this document, whether print or electronic, including but not limited to any current or future file format.

The Eight-Fold way June 4, 2007

Physics 213 General Physics I

However, the requirements (i) that the full copyright notice be included; (ii) that the disclaimer of warranty be included; and (iii) that the license text or its URI be included; as detailed in sections 4(a) and 4(d) of the license, are hereby waived by the author provided all of the following conditions are met:

1. the copies are distributed exclusively to students enrolled in a course—which is taught by the licensee—at a secondary or post-secondary school;

2. the copies are distributed in hard copy or binary-only form (e.g., PDF); if distributed in binary form via a computer network, the URI of the copy must be restricted to students enrolled in the course;

3. the author is credited in the following or some substantially similar form: “Written by Itai Seggev” for copies of the Work and “Based on a Work by Itai Seggev” for derivatives of the Work.

For details of the license see http://www.creativecommons.org/licenses/by-nc-sa/3.0/.